# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

### v4

@@ 1 Introduction
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated. Recent work towards brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive (Moses et al., 2021; Willett et al., 2023) or non-invasive brain recordings (Tang et al., 2023), mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain, especially in magnetoencephalography (MEG) data. Current speech decoding models rarely train on multiple subjects, combine datasets, or utilise data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations holds the promise of training models on collective, internet-scale data.
@@ 3.2 Pretext Tasks
To learn representations that encode phase differences between brain areas, this task applies a discrete uniform random phase shift $\phi\in\{0,\frac{\pi}{8},\frac{\pi}{4},\frac{3\pi}{8},\frac{\pi}{2},\frac{5\pi}{8},\frac{3\pi}{4},\frac{7\pi}{8}\}$ to a uniformly randomly selected proportion $\rho$ of the sensors. Applying this shift to random sensors is critical since sensors are placed in different positions, capturing different regions of the brain. Uniform random selection ensures differences between any two regions of the brain are represented. The objective of this task is to predict the phase shift. This leads to a similar loss
@@
Thus, this pretext task focuses on learning representations that encode relative sensor amplitude differences. Similar to the phase shift task, we select a random proportion of the sensors $\rho$ and apply a discrete random amplitude scaling coefficient $A\in[-2,2]$, discretised into 16 scaling factors, to the signal. The objective is to predict the scaling factor, leading to the loss
@@
where $w_{i}$ is a constant coefficient for each self-supervised loss.
@@ 4.2 Learning Generalisable Representations Using Pretext Tasks
1Yang et al. (2023) 2Wang et al. (2023)
Table 2: Our approach outperforms all baselines and matches surgical decoding. In the linear-only case, we train a supervised linear classifier on the input MEG signals. For ours and BrainBERT, we train a linear layer on top of a backbone pre-trained on CamCAN, with the rest of the model frozen. For BIOT, we use their pre-trained weights. In the no pre-training baseline, the backbone uses randomly initialised and frozen weights. In the surgical context, we quote the relevant result from (Wang et al., 2023, Table 2). With all pretext tasks, their losses are weighted equally.
@@
Our last baseline is BIOT, which we were unable to train successfully beyond chance level using MEG data. Instead, in (8), we leverage publicly released weights pre-trained with over fifty thousand hours of EEG. However, BIOT still falls short of our pre-training method for reasons which we believe are similar to BrainBERT—its objective does not leverage neuroscientific understanding of speech processing.
@@ 4.5 Aggregating Unlabelled MEG Datasets
Table 3: Aggregating unlabelled datasets outperforms single studies. For the first time with MEG data, we show that unlabelled pre-training data from multiple studies with different hardware profiles can be aggregated while continuing to gain the benefits of scaling. Combining data leads to a statistically significant $(p<0.05)$ improvement.
@@ Appendix A Experiment Details
BrainBERT adaptation. As BrainBERT is designed for single-electrode representations, for a fair comparison, to take into account all sensors, we ensured our linear classifier is applied to a concatenated embedding over all sensors. This is a large vector and leads to very computationally expensive training and is the reason we had to reduce the pre-training and downstream data in part C of Table 2.
@@ Appendix B Hyperparameters
Why $\rho$? The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search. We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors. Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset (Hamilton et al., 2018). In contrast, a larger $\rho$ for phase shift prediction could lead to representations that better discriminate neural synchrony information which is distributed across the brain rather than localised. As a result, a large proportion of the sensors in a MEG scanner should encode information about this feature.

### v5

@@ 1 Introduction
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet been fully assimilated. Recent work towards brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive (Moses et al., 2021; Willett et al., 2023) or non-invasive brain recordings (Tang et al., 2023), mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain, especially in magnetoencephalography (MEG) data. Current speech decoding models rarely train on multiple subjects, combine datasets, or utilise data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations holds the promise of training models on collective, internet-scale data. https://pnpl.robots.ox.ac.uk/bbl
@@ 3.2 Pretext Tasks
To learn representations that encode phase differences between brain areas, this task applies a discrete uniform random phase shift $\phi\in\{0,\frac{\pi}{8},\frac{\pi}{4},\frac{3\pi}{8},\frac{\pi}{2},\frac{5\pi}{8},\frac{3\pi}{4},\frac{7\pi}{8}\}$ to a uniformly randomly selected proportion $\rho\in[0,0.5]$ of the sensors. Applying this shift to random sensors is critical since sensors are placed in different positions, capturing different regions of the brain. Uniform random selection ensures differences between any two regions of the brain are represented. The objective of this task is to predict the phase shift. This leads to a similar loss
@@
Thus, this pretext task focuses on learning representations that encode relative sensor amplitude differences. Similar to the phase shift task, we select a random proportion of the sensors $\rho\in[0,0.5]$ and apply a discrete random amplitude scaling coefficient $A\in[-2,2]$, discretised into 16 scaling factors, to the signal. The objective is to predict the scaling factor, leading to the loss
@@
where $w_{i}$ is a constant coefficient for each loss.
@@ 4.2 Learning Generalisable Representations Using Pretext Tasks
1Yang et al. (2023) 2Wang et al. (2023) 3Wang et al. (2024)
Table 2: Our approach surpasses baselines in speech detection by up to 27% and matches surgical decoding. For linear, we train a supervised linear classifier on the MEG signals. For ours and BrainBERT, we train a linear layer on top of a backbone pre-trained on CamCAN, with the rest of the model frozen. For BIOT, we use their pre-trained weights. In the no pre-training baseline, the backbone uses randomly initialised and frozen weights. In the surgical context, we quote the result from Wang et al. (2023, Table 2). With all pretext tasks, losses are weighted equally.
@@
Our last baselines are BIOT (8) and EEGPT (10). We leverage publicly released weights pre-trained with thousands of hours of EEG. However, both still fall short of our pre-training method for reasons which we believe are similar to BrainBERT—their objectives do not leverage neuroscientific understanding of speech processing.
@@ 4.5 Aggregating Unlabelled MEG Datasets
Table 3: Aggregating unlabelled datasets outperforms single studies in speech detection. For the first time with MEG, we show that unlabelled pre-training data from multiple studies with different hardware profiles can be aggregated while gaining the benefits of scaling. Combining data leads to a significant $(p<0.05)$ improvement.
@@ 5 Conclusion
Acknowledgements
We would like to thank Botos Csaba for many early insightful discussions which helped shaped the direction of this work. In alphabetical order, thanks also to Mats W.J. van Es for technical assistance with the OSL library, Yonatan Gideoni for advice on data splits, Minqi Jiang for an encouraging conversation on scaling unsupervised representation learning, Brian Liu for technical contributions which did not reach the final paper, and Miran Özdogan for reviewing a draft of this work. Finally, we thank the rest of the PNPL group for their continued and unwavering support.
The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. http://dx.doi.org/10.5281/zenodo.22558.
[…]
@@ Appendix A Experiment Details
Adapting models for more sensors. As BrainBERT is designed for single-electrode representations, for a fair comparison, to take into account all sensors, we ensured our linear classifier is applied to a concatenated embedding over all sensors. This is a large vector and leads to very computationally expensive training and is the reason we had to reduce the pre-training and downstream data in part C of Table 2. We similarly concatenate sets of 19 sensors when evaluating EEGPT.
@@ Appendix B Hyperparameters
[…]